This number problem is a key to me understand several more:
Find two numbers whose difference is 12 if 4 times the smaller number is 7 less then 3 times the larger number.
I've gotten this far but i don't think it's right
small = x
large = x + 12
4x = 3(x + 12) - 7

Answer by ankor@dixie-net.com(22740)

Find two numbers whose difference is 12 if 4 times the smaller number is 7 less then 3 times the larger number.
I've gotten this far but i don't think it's right
small = x
large = x + 12
4x = 3(x + 12) - 7
this is right, continue
4x = 3x + 36 - 7
4x - 3x = 29
x = 29 is the smaller
and
29 + 12 = 41 is the larger